OpenStudy (anonymous):

X+Y+Z=1 -x+0.4y+0.6z=0 0.3x-y+0.7z=0 0.1x+0.9y-z=0

OpenStudy (anonymous):

Please help.. my answer was x=5/72 y=35/72 z=4/9 but it was wrong..